zoella:

if a2k= ak-1 × ak+1 2 den Evil is a1× an= a2× an-1=…= ak× an-k+1 3. im confused

YouMyPlug:

When making a post, or making comments on anywhere in the website, refrain from using curse words/profanity. It is against the rules and could unfortunately lead to an account suspension. With that being said, I hope you get help with your work.

kamariana:

A1 x An = a2 x an-1 =…= Ak x An - k + 1

fwval:

it should give you the same result for both sides

fwval:

correct me if im wrong

PureSoulless:

If I'm going to have to go through this in the future, might as well give up right now.

sllo:

Alright, so imagine your sequence is this weirdly balanced, overly organized family where every even-numbered sibling is just the mathematical lovechild of their two nearest odd-numbered siblings, but divided by 2 for fairness because, you know, equity or whatever:, . Now, just when you think, “Okay, that’s already kind of weird,” math hits you with another plot twist: if you take the first and last term and multiply them, then do the same with the second and second-last, and keep going, every single pair gives you the same result. Like some creepy secret society where everyone swears an oath to always multiply into the same number. At first, you’re like, “, what is happening?” But then it clicks. your sequence is basically mirroring itself in a multiplication dance, keeping everything weirdly stable, like a perfectly choreographed math conspiracy. it’s just a fancy way of saying the sequence is multiplying in a super-organized, mirror-like way. First and last, second and second-last, they all shake hands (or numbers) and agree to give you the same product, no matter what. At first, it feels like some cursed math prophecy, but once you get it, it’s just a predictable, well-behaved pattern.

surjithayer:

\[(a _{k})^2=a _{k-1}\times a _{k+1}\] \[\frac{ a _{_{k}} }{ a _{k-1} }=\frac{ a _{k+1} }{ a _{k} }\] it means the terms of a sequence are in G.P. \[\because ~common~ratio~of~successive~terms~is~same.\]